OpenStudy (anonymous):

Help please. 9|9-8x| =2x+3

OpenStudy (anonymous):

You must divide the absolute value function into two. The first one would be the positive value of the function 9(9-8x)=2x+3 81-72x=2x+3 Then solve for x The second one would be the negative value of the function -9(9-8x)=2x+3 -81+72x=2x+3 Then solve for x